In a performance test, each of two cars takes 8.9 s to accelerate from rest to 28 m/s. Car A has a mass of 1429 kg, and car B has a mass of 1889 kg. Find the net average force that acts on (a) car A and (b) car B during the test.
Answer #1

Solution :-

initial velocity = 28 m/s

t =8.9 sec

average acceleration = (final velocity - initial velocity)/t
=> 28/8.9 = 3.146 m/s^2

(a) average force acting on car A = 1429*3.146 = 4495.6 N
(b) average force acting on car B = 1889*3.146 = 5942.8 N